OpenStudy (anonymous):

factor: 27x^3y^6+8p^3

hartnn (hartnn):

can u write 27x^3y^6 as (....)^3 what is .... ?

OpenStudy (anonymous):

what?

hartnn (hartnn):

27 is cube of what number ?

OpenStudy (anonymous):

3

hartnn (hartnn):

yes, so u can write \(\huge 27x^3y^6 = (3xy^2)^3\) got this ?

OpenStudy (anonymous):

no

hartnn (hartnn):

(3xy^2)^3 = 3^3 x^3 (y^2)^3 = 27 x^3 y^6 now ?

OpenStudy (anonymous):

the answer is \[(3xy^2+2p)(9x^y^4-6xy^2p+4p^2)\]

OpenStudy (anonymous):

just dont knw how to get it

hartnn (hartnn):

\(\huge 27x^3y^6 = (3xy^2)^3 \\ \huge 8p^3= (2p)^3\) now use \(\huge a^3+b^3 = (a+b)(a^2-ab+b^2)\)

hartnn (hartnn):

taking a= (3xy^2) and b= 2p

jimthompson5910 (jim_thompson5910):

First rewrite as a sum of cubes 27x^3y^6+8p^3 (3xy^2)^3+8p^3 (3xy^2)^3+(2p)^3 Then use the sum of cubes factoring formula a^3 + b^3 = (a+b)(a^2 - ab + b^2) to get In this case, a = 3xy^2 and b = 2p (3xy^2 + 2p)( (3xy^2)^2 - (3xy^2)*(2p) + (2p)^2 ) (3xy^2 + 2p)( 9x^2y^4 - 6xy^2p + 4p^2 )

OpenStudy (anonymous):

thank you for explaining lol

jimthompson5910 (jim_thompson5910):

yw